OpenStudy (anonymous):

The length of a rectangle is 2 inches less than 3 times the number of inches in its width. If the perimeter of the rectangle is 28 inches, what is the width and length of the rectangle? {Width_________________ a0in., Length_______________ a1in.}

OpenStudy (campbell_st):

let the width be w and then the length is l = 3w - 2 so P=2(l + w) or 28 = 2(w + 3w -2) now solve for w... hope it helps

OpenStudy (anonymous):

so is W=4

OpenStudy (campbell_st):

yep... the width is w and the length is 3*4 - 2

OpenStudy (campbell_st):

lol..opps width is 4

OpenStudy (anonymous):

wait what?

OpenStudy (campbell_st):

well the width is 4.... thats correct and you know the length l = 3w - 2 so make the substitution to find the length l = 3x4 - 2 l = 10

OpenStudy (anonymous):

okay thank you